2 Moles of CO2 gas contain the same number of atoms as _ moles of Co 

Grade:10

1 Answers

Debarati
2 years ago
No. of atoms in 1 mole of CO2= 3 (1 Carbon and 2 Oxygen)
No. of atoms in 2 mole of CO2= 3*2= 6
No. of atoms in 1 mole of CO= 2
No. of CO mole containing 6 atoms= 6/2= 3

I hope you understood.